OpenStudy (anonymous):

Adding and Subtracting Fractions and Mixed Numbers – Find the Exact Solution 1. 5/7 + 8/9 = 2. 4/5 – 1/2 = 3. 2 5/8 + 2/9 = 4. 9 5/6 – 3 1/3 = 5. 17/19 + 4 = show work

OpenStudy (anonymous):

?

OpenStudy (anonymous):

STOP SPAMMING

OpenStudy (anonymous):

YOU SAID THAT LIKE 3 OR 4 TIMES?!

OpenStudy (anonymous):

no ones coming in and helping me

OpenStudy (anonymous):

your just annoying them.. post those questions in yahoo answer then..

OpenStudy (anonymous):

I'll help you... but you can post it once and ask on the chat if anyone will help you

OpenStudy (anonymous):

1. 5/7 + 8/9 = Make the two equivalent \[{5\over7}+{8\over9}\]\[{5\times9\over7\times9}+{8\times7\over7\times9}\] \[{45\over63}+{56\over63}\]\[101\over63\]

OpenStudy (anonymous):

or 1\[1 {38\over63}\]

OpenStudy (anonymous):

See if you can do the rest by yourself?

OpenStudy (anonymous):

PLEASE FOUR MORE??

OpenStudy (anonymous):

I know there are four more. Try doing one on your own, and see if you can solve it. Then tell me you're answer, and I'll see if it's right.

OpenStudy (anonymous):

i cant i tried?? please

OpenStudy (anonymous):

I'll do one more... but not all of it. The steps in arriving at the solution is the same for all of them...

OpenStudy (anonymous):

Why did you delete it?

OpenStudy (anonymous):

sorry. 2. 4/5 – 1/2 = \[{4\times2\over5\times2}+{(-1)\times5\over5\times2}\] See if you can get an answer from here

OpenStudy (anonymous):

Is thisa example?

OpenStudy (anonymous):

No. this is half the step to the solution of the 2nd question you asked. See if you can get the answer from there.

OpenStudy (anonymous):

it will = 3/10

OpenStudy (anonymous):

Yes, correct :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Now, do the same with the rest!